# Multi-Modal Language Models as Text-to-Image Model Evaluators

### v1

@@ 1 Introduction
In our approach, we first focus on validating the capabilities of off-the-shelf MLLMs to generate questions and score prompt-generation consistency of T2I models, showing that a single model can outperform existing human judgment correlation metrics on the static TIFA160 benchmark (Hu et al., 2023). This contrasts with previous work that either relied on proprietary GPT-4 variants (Zhang et al., 2023) or found open-source MLLMs’ judgments ineffective and uncorrelated with human judgments (Ku et al., 2024; Jiang et al., 2024). Second, we leverage MLLMs to not only score but also generate prompts of increasing complexity to evaluate T2I models. We use this MLLM-based approach to compare the prompt-generation consistency of 8 state-of-the-art T2I models and find that the rankings from our approach are comparable to those reported by static benchmarks, while only using a fraction ($\sim$1%) of prompts. We also show that MT2IE is easily generalizable to evaluate aspects of image generation beyond prompt-generation consistency, with results for aesthetic quality presented in Appendix appendix A. Finally, we investigate MLLMs as evaluator agents for prompt-generation consistency by enabling the interaction between the MLLM and the T2I model.
@@ 3 Experimental Overview
We demonstrate the effectiveness of using MLLMs as evaluator agents for assessing T2I prompt-generation consistency. In Multimodal Text-to-Image Eval (MT2IE), MLLM interacts with the T2I model by proposing text prompts and scoring resulting generations. We show that our interactive approach enables MLLMs to effectively benchmark T2I models in two distinct settings: either by systematically generating prompts of increasing complexity or by dynamically adapting prompts based on T2I model performance. In the following sections, we focus on evaluating image-text consistency as it directly impacts the overall visual quality of generated images (Astolfi et al., 2024; Xu et al., 2023; Kirstain et al., 2023). Additionally, in Appendix appendix A we show that MT2IE is an easily generalizable framework, capable of evaluating other aspects of T2I models such as image aesthetics. We use MT2IE to evaluate a total of 8 recent T2I models including latent diffusion model (LDM) v2.1 (Rombach et al., 2022), LDM v3 (Esser et al., 2024), LDM XL (Podell et al., 2023), LDM XL-Turbo (Sauer et al., 2025), Flux (Labs, 2024), a cascaded pixel-based diffusion model (CDM) (Raffel et al., 2020), Playground 2.5 (Li et al., 2024b) and Pixart-$\alpha$ (Chen et al., 2023). We provide an outline of our experiments and elaborate in subsequent sections.
@@
We validate that open-sourced MLLMs are capable of scoring image-text consistency on existing benchmarks without the usage of other models or fine-tuning. Existing works that use MLLMs for end-to-end evaluation either do not use existing benchmarks or fine-tune open-sourced MLLMs for the task (Tan et al., 2024; Ku et al., 2024). Existing GQA-based benchmarks rely on a system of models for question generation and scoring; typically an LLM is used to generate questions based on a curated text prompt, the generated questions are validated by an additional LLM or question-answering model, then finally a VQA model scores the image using the generated questions (Cho et al., 2024; Yarom et al., 2023; Hu et al., 2023; Gupta and Kembhavi, 2022). We validate that MLLMs are capable of emulating GQA-based benchmarks by generating questions about a given prompt and scoring the generated questions given the corresponding image. We evaluate using TIFA160 (Hu et al., 2023), and show that a single open-sourced MLLM can recreate this benchmark and score questions with better correlation to human judgment. We detail this experiment in Section section 4.
@@
After validating MLLMs’ abilities to generate and score prompt consistency questions in Experiment 1, we then introduce MT2IE , demonstrating how MLLMs can iteratively generate a full T2I benchmark and judge model performance. In particular, we show that MLLM-generated benchmarks preserve relative model rankings computed by human annotators in prior work. We also show that our method is efficient, accurately ranking models with significantly fewer prompts than existing benchmarks of handwritten prompts. We detail these experiments in Section section 5.
@@
Beyond just generating benchmarks, we demonstrate that MT2IE can act as an evaluator agent, adapting to model-specific performance. Specifically, we show that MLLMs can adjust the complexity of prompts based on model performance on previous prompts. This enables each individual T2I model to have bespoke prompts tailored to its abilities, discussed in Section section 6.
@@ 4.1 Generated Question Answering Approach
GQA-based approaches use an LLM to generate questions based on prompts then score generated questions and images using a VQA model. The performance of the VQA model being used is crucial as VQA accuracy has direct correlation with human judgment (Hu et al., 2023). For this reason, we first confirm that MLLMs have strong VQA performance on existing GQA benchmarks by using the TIFA160 LLM-generated question set. We compare the MLLMs’ VQA accuracy to the 2 best VQA models from Hu et al. (2023); results are in Figure figure 9 (Appendix appendix C) and show that all 3 MLLMs outperform the VQA models used in existing GQA benchmarks.
@@
With this validation, we now investigate if a single MLLM can effectively generate and score questions given a text prompt and generated image. This constitutes a benchmark generation task, where the MLLM is responsible for creating new questions and scoring them. By consolidating these tasks to one model, this approach eliminates the resource-intensive reliance on external, often proprietary LLMs. We use all 3 MLLMs to generate and score questions about all prompts in TIFA160. The system prompts used to generate questions is shown in Appendix appendix C.
To assess quality of the generated questions, we follow previous work and measure the correlation between the VQA accuracy (from LLM-generated questions from existing benchmarks (Cho et al., 2024; Hu et al., 2023; Gupta and Kembhavi, 2022) or MLLM-generated questions in our work) and human ratings of image-text consistency using a 1-5 Likert scale. We use images and human ratings from prior work (Cho et al., 2024) on 5 different T2I models: LDM1.1 (Rombach et al., 2022), LDM1.5 (Rombach et al., 2022), LDM2.1 (Rombach et al., 2022), minDALL-E (Saehoon et al., 2021), and VQ-Diffusion (Gu et al., 2022). We show the Spearman’s $\rho$ and Kendall’s $\tau$ in Table table 1.
Notably, every MLLM achieves higher correlation with human annotations than existing GQA-based approaches. These results show that open-sourced MLLMs can effectively execute all tasks performed by separate models in existing GQA benchmarks. We additionally compare the GPT3-generated questions in TIFA160 and the MLLM-generated questions from Llava-v1.6-34b, the LLM with the highest human correlations. We analyze question statistics and BLEU score (Papineni et al., 2002) (a commonly used NLP metric to compare two sets of text) and find that entities in both sets of questions are very similar. Detailed results on generated question set verification are in Appendix appendix G.
@@ 4.2 VQAScore Approach
VQAScore (Lin et al., 2024) computes the likelihood of the model generating “yes" when asked “Does this figure show {text prompt}. Please answer yes or no.". VQAScore has been shown to be more reliable than GQA-based approaches and alignment metrics like CLIPScore, likely due to CLIPScore’s challenges with object composition  (Lin et al., 2024; Tong et al., 2024; Thrush et al., 2022a; Yuksekgonul et al., 2023), and the tendency of GQA-based approaches to exhibit a “yes-bias" in question answering tasks (Ross et al., 2024). We compare the correlations with human ratings on TIFA160 between VQAScore calculated using the 3 aforementioned MLLMs, VQAScore calculated with the CLIP-FlanT5 model (which was fine-tuned for SOTA VQAScore calculation (Lin et al., 2024)), and CLIPScore. We use the same human ratings as in Section section 4.1; results are in Table table 2. All MLLMs outperform CLIPScore in human correlation, but only Llava-v1.6-34b achieves higher correlation with human ratings than CLIP-FlanT5.
We find that open-sourced MLLMs can replace systems of LLM and VQA models commonly used in existing QGA benchmarks. Llava-v1.6-34b achieved the best correlation with human ratings when used to compute VQAScore as prompt consistency scores; so we use Llava-v1.6-34b with VQAScore as our MLLM evaluator for further experiments. Experiments with question generation and answering for prompt consistency scoring are in Appendix appendix C. This validation lays the framework for our experiments, demonstrating that open-sourced MLLMs can be used for end-to-end T2I evaluation.
@@ 5 Experiment 2: MLLMs for Benchmark Generation
In Section section 4, we showed that MLLMs are able to reproduce static benchmarks using both question-answering and prompt likelihood approaches. We now introduce our approach, the MT2IE framework, and show that it can be used to iteratively generate new, efficient benchmarks for image-text consistency. With each iteration, the generated prompts increase in complexity and thus challenge the T2I models in progressively more difficult settings. This involves first providing a starting prompt, which we generate in our experiments but can be passed in by a user, that we refer to as a seed prompt; then expanding on the seed prompt to generate increasingly more difficult and complex prompts. MT2IE-generated benchmarks are effective in that they match prior model rankings from benchmarks of handwritten prompts, while also being efficient in that they require far fewer examples than existing benchmarks. Additionally, MT2IE-generated benchmarks are dynamic, potentially addressing the dataset leakage and saturation issues of static benchmarks.
@@
Our approach iterates between two stages, shown on the right side of Figure figure 1. First, during iterative prompt generation, the MLLM is given an existing prompt $T_{i-1}$ and generates a new prompt $T_{i}$ that builds on $T_{i-1}$ to become more difficult. $T_{i}$ contains an additional object, spatial relationship, or object attribute compared to $T_{i-1}$, incrementally increasing the complexity of the prompt every iteration; see Figure figure 2 for an example. The seed prompt $T_{0}$ can be user-written or generated. The T2I model generates an image $I_{i}$ based on the prompt $T_{i}$. Next, during image scoring, the MLLM scores $I_{i}$ based on its faithfulness to prompt $T_{i}$. A full run consists of $N$ iterations of both iterative prompt generation and image scoring stages, with each iteration yielding a score $s_{i}$. We perform 756 full runs, with $N=5$ iterations per-run, resulting in a total of 3,780 scores from all runs. For these runs we evaluate a subset of 4 T2I models: CDM, LDM v2.1, LDM XL and LDM XL-Turbo. Our framework is flexible and supports any scoring metric during the image scoring stage; in this paper we use VQAScore (Lin et al., 2024), as it achieved the highest correlations with human judgment (see Experiment 1 in Section section 4.2).
Figure figure 3(a) shows average scores for each prompt iteration. We find that scores for all models decrease as prompt iterations increase, indicating that MT2IE effectively generates prompts that are progressively more challenging. To support this, we also analyze the change in linguistic properties of prompts across iterations in Figure figure 3(b); we use the Flesh-Kincaid score (Flesch, 1948) to compute semantic difficulty. As MT2IE-generated prompts get more difficult, they still maintain coherence and grammaticality. In general, prompts increase in their word count and semantic difficulty as prompt iterations occur. See an example progression of a 5-iteration run in Figure figure 2.
@@
After establishing that MT2IE generates progressively challenging prompt consistency benchmarks and achieves higher correlations in scoring generations, we next investigate how efficient MT2IE-generated benchmarks are compared to static benchmarks. Concretely, we ask: can we achieve the same evaluation results, i.e.,  same model rankings, while requiring fewer text prompts? We start with 4 generated seed prompts across 4 topic categories: household scenes (foods, household items, or furniture), descriptions of people, scenes with animals, and location descriptions. Each seed prompt is moderately complex (2-3 compositional concepts per prompt) and undergoes 4 difficulty progression iterations ($T_{0}\rightarrow T_{4}$), resulting in 20 total evaluation prompts in the generated benchmark (4 seeds $\times$ 5 prompts). Empirically we found that 5 prompts, resulting from 4 iterations, provided a sufficient amount of prompt difficulty (see Appendix appendix F for ablations on number of iterations). We refer to this as MT2IEbench20.
Table 3: Rank correlations between T2I model rankings on full GenAIBench (1600 prompts) versus only 20 prompts using our method versus all other evaluation methods (also see Figure figure 4). MT2IE produces notably high correlations rankings.
@@
We rank the 8 T2I models listed in Section section 3 by (i) average score on 20 generated prompts from MT2IEbench20 and (ii) VQAScore evaluated on 1600 human-written prompts from GenAIBench (Li et al., 2024a). Next, for a direct comparison we randomly sample 20 prompts from GenAIBench, ensuring uniform sampling across all prompt categories, and refer to this as GenAIBench20. We rank the T2I models on GenAIBench20 by (iii) VQAScore (iv) CLIPScore (v) VIEScore.We run MT2IE-benchmark generation over 5 random seeds to ensure results are statistically significant (model score mean and standard deviations in Figure figure 7 and Figure figure 11, Appendix appendix C). Because variance across 5 seeds was extremely high for all metrics ran on GenAIBench20, all GenAIBench20 results are reported over 10 seeds. The relative rankings and scores for each evaluation method are shown in Figure figure 4.

### v2

@@ 1 Introduction
In our approach, we first focus on validating the capabilities of off-the-shelf MLLMs to generate questions and score prompt-generation consistency of T2I models, showing that a single model can outperform existing human judgment correlation metrics on the static TIFA160 benchmark (Hu et al., 2023). This contrasts with previous work that either relied on proprietary GPT-4 variants (Zhang et al., 2023) or found open-source MLLMs’ judgments ineffective and uncorrelated with human judgments (Ku et al., 2024; Jiang et al., 2024). Second, we leverage MLLMs to not only score but also generate prompts of increasing complexity to evaluate T2I models. We use this MLLM-based approach to compare the prompt-generation consistency of 8 state-of-the-art T2I models and find that the rankings from our approach are comparable to those reported by static benchmarks, while only using a fraction ($\sim$1%) of prompts. We also show that MT2IE is easily generalizable to evaluate aspects of image generation beyond prompt-generation consistency, with results for aesthetic quality presented in Appendix A. Finally, we investigate MLLMs as evaluator agents for prompt-generation consistency by enabling the interaction between the MLLM and the T2I model.
@@ 3 Experimental Overview
We demonstrate the effectiveness of using MLLMs as evaluator agents for assessing T2I prompt-generation consistency. In Multimodal Text-to-Image Eval (MT2IE), MLLM interacts with the T2I model by proposing text prompts and scoring resulting generations. We show that our interactive approach enables MLLMs to effectively benchmark T2I models in two distinct settings: either by systematically generating prompts of increasing complexity or by dynamically adapting prompts based on T2I model performance. In the following sections, we focus on evaluating image-text consistency as it directly impacts the overall visual quality of generated images (Astolfi et al., 2024; Xu et al., 2023; Kirstain et al., 2023). Additionally, in Appendix A we show that MT2IE is an easily generalizable framework, capable of evaluating other aspects of T2I models such as image aesthetics. We use MT2IE to evaluate a total of 8 recent T2I models including latent diffusion model (LDM) v2.1 (Rombach et al., 2022), LDM v3 (Esser et al., 2024), LDM XL (Podell et al., 2023), LDM XL-Turbo (Sauer et al., 2025), Flux (Labs, 2024), a cascaded pixel-based diffusion model (CDM) (Raffel et al., 2020), Playground 2.5 (Li et al., 2024b) and Pixart-$\alpha$ (Chen et al., 2023). We provide an outline of our experiments and elaborate in subsequent sections.
@@
We validate that open-sourced MLLMs are capable of scoring image-text consistency on existing benchmarks without the usage of other models or fine-tuning. Existing works that use MLLMs for end-to-end evaluation either do not use existing benchmarks or fine-tune open-sourced MLLMs for the task (Tan et al., 2024; Ku et al., 2024). Existing GQA-based benchmarks rely on a system of models for question generation and scoring; typically an LLM is used to generate questions based on a curated text prompt, the generated questions are validated by an additional LLM or question-answering model, then finally a VQA model scores the image using the generated questions (Cho et al., 2024; Yarom et al., 2023; Hu et al., 2023; Gupta and Kembhavi, 2022). We validate that MLLMs are capable of emulating GQA-based benchmarks by generating questions about a given prompt and scoring the generated questions given the corresponding image. We evaluate using TIFA160 (Hu et al., 2023), and show that a single open-sourced MLLM can recreate this benchmark and score questions with better correlation to human judgment. We detail this experiment in Section 4.
@@
After validating MLLMs’ abilities to generate and score prompt consistency questions in Experiment 1, we then introduce MT2IE , demonstrating how MLLMs can iteratively generate a full T2I benchmark and judge model performance. In particular, we show that MLLM-generated benchmarks preserve relative model rankings computed by human annotators in prior work. We also show that our method is efficient, accurately ranking models with significantly fewer prompts than existing benchmarks of handwritten prompts. We detail these experiments in Section 5.
@@
Beyond just generating benchmarks, we demonstrate that MT2IE can act as an evaluator agent, adapting to model-specific performance. Specifically, we show that MLLMs can adjust the complexity of prompts based on model performance on previous prompts. This enables each individual T2I model to have bespoke prompts tailored to its abilities, discussed in Section 6.
@@ 4.1 Generated Question Answering Approach
GQA-based approaches use an LLM to generate questions based on prompts then score generated questions and images using a VQA model. The performance of the VQA model being used is crucial as VQA accuracy has direct correlation with human judgment (Hu et al., 2023). For this reason, we first confirm that MLLMs have strong VQA performance on existing GQA benchmarks by using the TIFA160 LLM-generated question set. We compare the MLLMs’ VQA accuracy to the 2 best VQA models from Hu et al. (2023); results are in Figure 9 (Appendix C) and show that all 3 MLLMs outperform the VQA models used in existing GQA benchmarks.
@@
With this validation, we now investigate if a single MLLM can effectively generate and score questions given a text prompt and generated image. This constitutes a benchmark generation task, where the MLLM is responsible for creating new questions and scoring them. By consolidating these tasks to one model, this approach eliminates the resource-intensive reliance on external, often proprietary LLMs. We use all 3 MLLMs to generate and score questions about all prompts in TIFA160. The system prompts used to generate questions is shown in Appendix C.
To assess quality of the generated questions, we follow previous work and measure the correlation between the VQA accuracy (from LLM-generated questions from existing benchmarks (Cho et al., 2024; Hu et al., 2023; Gupta and Kembhavi, 2022) or MLLM-generated questions in our work) and human ratings of image-text consistency using a 1-5 Likert scale. We use images and human ratings from prior work (Cho et al., 2024) on 5 different T2I models: LDM1.1 (Rombach et al., 2022), LDM1.5 (Rombach et al., 2022), LDM2.1 (Rombach et al., 2022), minDALL-E (Saehoon et al., 2021), and VQ-Diffusion (Gu et al., 2022). We show the Spearman’s $\rho$ and Kendall’s $\tau$ in Table 1.
Notably, every MLLM achieves higher correlation with human annotations than existing GQA-based approaches. These results show that open-sourced MLLMs can effectively execute all tasks performed by separate models in existing GQA benchmarks. We additionally compare the GPT3-generated questions in TIFA160 and the MLLM-generated questions from Llava-v1.6-34b, the LLM with the highest human correlations. We analyze question statistics and BLEU score (Papineni et al., 2002) (a commonly used NLP metric to compare two sets of text) and find that entities in both sets of questions are very similar. Detailed results on generated question set verification are in Appendix G.
@@ 4.2 VQAScore Approach
VQAScore (Lin et al., 2024) computes the likelihood of the model generating “yes" when asked “Does this figure show {text prompt}. Please answer yes or no.". VQAScore has been shown to be more reliable than GQA-based approaches and alignment metrics like CLIPScore, likely due to CLIPScore’s challenges with object composition  (Lin et al., 2024; Tong et al., 2024; Thrush et al., 2022a; Yuksekgonul et al., 2023), and the tendency of GQA-based approaches to exhibit a “yes-bias" in question answering tasks (Ross et al., 2024). We compare the correlations with human ratings on TIFA160 between VQAScore calculated using the 3 aforementioned MLLMs, VQAScore calculated with the CLIP-FlanT5 model (which was fine-tuned for SOTA VQAScore calculation (Lin et al., 2024)), and CLIPScore. We use the same human ratings as in Section 4.1; results are in Table 2. All MLLMs outperform CLIPScore in human correlation, but only Llava-v1.6-34b achieves higher correlation with human ratings than CLIP-FlanT5.
We find that open-sourced MLLMs can replace systems of LLM and VQA models commonly used in existing QGA benchmarks. Llava-v1.6-34b achieved the best correlation with human ratings when used to compute VQAScore as prompt consistency scores; so we use Llava-v1.6-34b with VQAScore as our MLLM evaluator for further experiments. Experiments with question generation and answering for prompt consistency scoring are in Appendix C. This validation lays the framework for our experiments, demonstrating that open-sourced MLLMs can be used for end-to-end T2I evaluation.
@@ 5 Experiment 2: MLLMs for Benchmark Generation
In Section 4, we showed that MLLMs are able to reproduce static benchmarks using both question-answering and prompt likelihood approaches. We now introduce our approach, the MT2IE framework, and show that it can be used to iteratively generate new, efficient benchmarks for image-text consistency. With each iteration, the generated prompts increase in complexity and thus challenge the T2I models in progressively more difficult settings. This involves first providing a starting prompt, which we generate in our experiments but can be passed in by a user, that we refer to as a seed prompt; then expanding on the seed prompt to generate increasingly more difficult and complex prompts. MT2IE-generated benchmarks are effective in that they match prior model rankings from benchmarks of handwritten prompts, while also being efficient in that they require far fewer examples than existing benchmarks. Additionally, MT2IE-generated benchmarks are dynamic, potentially addressing the dataset leakage and saturation issues of static benchmarks.
@@
Our approach iterates between two stages, shown on the right side of Figure 1. First, during iterative prompt generation, the MLLM is given an existing prompt $T_{i-1}$ and generates a new prompt $T_{i}$ that builds on $T_{i-1}$ to become more difficult. $T_{i}$ contains an additional object, spatial relationship, or object attribute compared to $T_{i-1}$, incrementally increasing the complexity of the prompt every iteration; see Figure 2 for an example. The seed prompt $T_{0}$ can be user-written or generated. The T2I model generates an image $I_{i}$ based on the prompt $T_{i}$. Next, during image scoring, the MLLM scores $I_{i}$ based on its faithfulness to prompt $T_{i}$. A full run consists of $N$ iterations of both iterative prompt generation and image scoring stages, with each iteration yielding a score $s_{i}$. We perform 756 full runs, with $N=5$ iterations per-run, resulting in a total of 3,780 scores from all runs. For these runs we evaluate a subset of 4 T2I models: CDM, LDM v2.1, LDM XL and LDM XL-Turbo. Our framework is flexible and supports any scoring metric during the image scoring stage; in this paper we use VQAScore (Lin et al., 2024), as it achieved the highest correlations with human judgment (see Experiment 1 in Section 4.2).
Figure 3(a) shows average scores for each prompt iteration. We find that scores for all models decrease as prompt iterations increase, indicating that MT2IE effectively generates prompts that are progressively more challenging. To support this, we also analyze the change in linguistic properties of prompts across iterations in Figure 3(b); we use the Flesh-Kincaid score (Flesch, 1948) to compute semantic difficulty. As MT2IE-generated prompts get more difficult, they still maintain coherence and grammaticality. In general, prompts increase in their word count and semantic difficulty as prompt iterations occur. See an example progression of a 5-iteration run in Figure 2.
@@
After establishing that MT2IE generates progressively challenging prompt consistency benchmarks and achieves higher correlations in scoring generations, we next investigate how efficient MT2IE-generated benchmarks are compared to static benchmarks. Concretely, we ask: can we achieve the same evaluation results, i.e.,  same model rankings, while requiring fewer text prompts? We start with 4 generated seed prompts across 4 topic categories: household scenes (foods, household items, or furniture), descriptions of people, scenes with animals, and location descriptions. Each seed prompt is moderately complex (2-3 compositional concepts per prompt) and undergoes 4 difficulty progression iterations ($T_{0}\rightarrow T_{4}$), resulting in 20 total evaluation prompts in the generated benchmark (4 seeds $\times$ 5 prompts). Empirically we found that 5 prompts, resulting from 4 iterations, provided a sufficient amount of prompt difficulty (see Appendix F for ablations on number of iterations). We refer to this as MT2IEbench20.
Table 3: Rank correlations between T2I model rankings on full GenAIBench (1600 prompts) versus only 20 prompts using our method versus all other evaluation methods (also see Figure 4). MT2IE produces notably high correlations rankings.
@@
We rank the 8 T2I models listed in Section 3 by (i) average score on 20 generated prompts from MT2IEbench20 and (ii) VQAScore evaluated on 1600 human-written prompts from GenAIBench (Li et al., 2024a). Next, for a direct comparison we randomly sample 20 prompts from GenAIBench, ensuring uniform sampling across all prompt categories, and refer to this as GenAIBench20. We rank the T2I models on GenAIBench20 by (iii) VQAScore (iv) CLIPScore (v) VIEScore.We run MT2IE-benchmark generation over 5 random seeds to ensure results are statistically significant (model score mean and standard deviations in Figure 7 and Figure 11, Appendix C). Because variance across 5 seeds was extremely high for all metrics ran on GenAIBench20, all GenAIBench20 results are reported over 10 seeds. The relative rankings and scores for each evaluation method are shown in Figure 4.